OpenStudy (anonymous):

3cos2(x)+2sin(x)+1=0 find all the values of x. a metal will be given

OpenStudy (anonymous):

replace \(3\cos^2(x)\) by \(3(1-\sin^2(x))\) and then you will have a quadratic equation in \(\sin(x)\)

OpenStudy (anonymous):

its not cos^2 its cos2

OpenStudy (anonymous):

\[3\cos 2(x)+2\sin(x)+1=0\]